A golf ball (mass 0.046kg) receives a linear impulse of 2.72kgm/s when it is struck with a driving iron. At what speed does the ball leave the tee?
1
Expert's answer
2020-01-29T14:08:04-0500

The speed is


"v=\frac{p}{m}"

"v=\frac{2.72}{0.046}=59.1\frac{m}{s}"